Lakeland Corporation earned $0.50 per share ten year ago. This year its earnings per share were $2.20. What was the growth rate in Lakeland’s earnings per share over the ten year period?
Answer #1

Let, EPS0 = $0.5

and EPS10 = $2.20

EPS10 = EPS0*(1+g)T

Where g = growth rate

T = time period

Therefore,

2.2 = 0.50*(1+g)10

(1+g)10 = 2.2/0.50

(1+g) = (4.4)1/10

g = 1.1597-1

g = 0.1597 or 15.97%

Therefore, growth rate over the ten year period was 15.97%.